OpenStudy (anonymous):

find the value of x so so that the line passing through (10,5) and (x,9) has a slope of -2 -8 8 2 -2

OpenStudy (owlcoffee):

So, all you can do is apply the equation of the slope, defined by: \[m_{a,b}=\frac{ y_2 - y_1 }{ x_2-x_1 }\] Which, in more simpler words mean: "the difference of the y coordinates divided by the difference of the x-coordinates". So, applying it for the points give, calling the coordinates (10,5) as point "a" and the coordinates (x,9) as point "b", then: \[m_{a,b}=\frac{ 9-5 }{ x-10 }\] But the desired slope is "-2", so therefore: \[-2=\frac{ 9-5 }{ x-10 }\] So all you have to do is solve for "x".

OpenStudy (anonymous):

I think it is -2